Does 7.8(x+3.1) equal 7 40/50+24 9/50. By the way I converted the decimals into fractions.
Answer by macston(5194)

.
Not necessarily.
If it does:
7.8(x+3.1)=7.8+24.18
7.8x+24.18=7.8+24.18
7.8x=7.8
x=1
They are only equal when x=1.